OpenStudy (anonymous):

The length of a rectangle is three times longer than its width. The perimeter of the rectangle is 16 feet. If x represents the width of the rectangle, which equation can be used to find its value?

OpenStudy (austinl):

The perimeter or a rectangle is, \(P=2l+2x\) And in this problem, \(l=3x\) So therefore, \(P=2(3x)+2x\) \(P=6x+2x=8x\)

OpenStudy (anonymous):

thank you @austinL

OpenStudy (austinl):

You're welcome! Can you solve it from there?

OpenStudy (anonymous):

yeah